A two digit number in base ten is equal to five times the sum of the digits. It is nine less than the number formed by interchanging the digits. Find the number.
Answer by lwsshak3(11628)

A two digit number in base ten is equal to five times the sum of the digits. It is nine less than the number formed by interchanging the digits. Find the number.
***
let u=units digit
let t=tens digit
..
10t+u=5(t+u)
10t+u=10u+t-9
..
10t+u=5t+5u
10t+u=10u+t-9
..
5t=4u
t=4u/5
..
10(4u/5)+u=10u+4u/5-9
8u+u=10u+4u/5-9
u+4u/5=9
5u+4u=45
9u=45
u=5
t=4u/5=4
number=45